OpenStudy (anonymous):

f the sequence is geometric, what is its common ratio? 512, 128, 32, 8,

OpenStudy (lalaly):

128/512 = 1/4

OpenStudy (anonymous):

what is the sequence is geometric, what is the common ratio? 48, 46, 44, 42

OpenStudy (lalaly):

46/48=23/24

OpenStudy (lalaly):

yu divide any term by the one that comes before it

OpenStudy (anonymous):

8x − 2 = y − 3 4x − y = -2

OpenStudy (lalaly):

8x-y=-1 4x-y=-2 subtract both equations u get 4x=1 so x=1/4 substitute x=1/4 in any equation yu get 2-y=-1 so y = 3